# Anymate: A Dataset and Baselines for Learning 3D Object Rigging

### v1

@@ Abstract.
ccs: Computing methodologies Animation
@@ 1. Introduction
With rapid advances of immersive media comes a growing demand for automated 3D content creation. 3D animation is essential nowadays across the entertainment industry, from game design and movie production to a wide range of AR/VR applications. Yet, producing realistic 3D animations remains one of the most labor-intensive tasks, often requiring enormous expert effort.
@@
The key insight is to leverage existing large-scale animated 3D assets handcrafted by expert artists. Large-scale 3D data has proven crucial in training high-quality 3D generation models (Shi et al., 2023; Liu et al., 2024). Notably, the Objaverse datasets (Deitke et al., 2022, 2024), aggregating over 10 million 3D assets from various public sources, have powered many state-of-the-art methods in recent years. However, prior efforts primarily used such assets for learning static 3D generation, neglecting the fact that many of these assets also come with valuable animation information skillfully crafted by the artists.
To train a learning-based model for 3D object rigging, we assemble a large-scale dataset of 230K rigged 3D assets curated from the Objaverse-XL Dataset (Deitke et al., 2024), each paired with artist-created rigging and skinning in a unified format, dubbed the Anymate Dataset. This dataset contains a wide spectrum of 3D objects, ranging from humanoid and animal characters to animated everyday objects like furniture and machines. The dataset is about 70 times larger than existing public rigging datasets, as summarized in Table 1.
@@ 3.2. Dataset Processing
We load these assets into Blender***https://www.blender.org/ and extract the mesh, armature (bones), and skinning weights. For training efficiency, we resample 8,192 points on the mesh, and recompute their skinning weights using barycentric interpolation based on the original annotated skinning weights at the vertices. Finally, for each asset, we export the original mesh in a unified .obj format, as well as the resampled point clouds and the corresponding bone locations and skinning weights.
@@ 4.2.1. Architectures
A regression-based model is perhaps the most intuitive approach for joint prediction, which simply regresses a set of 3D coordinates from the extracted shape features and can be directly supervised by the ground-truth joints during training. However, there are two unique features that make our joint prediction task different from typical regression problems: (1) the number of target joints varies across instances, and (2) the order of the joints is immaterial at this stageAlthough an order could technically be established based on the skeleton, inconsistencies may arise across instances created by different artists.. To address these challenges, we design a transformer-based architecture (Jaegle et al., 2021; Vaswani et al., 2017), which is permutation-equivariant and has proven to be highly scalable, followed by a clustering step (Ester et al., 1996).
@@
where $\overline{\alpha}_{t}$ signifies the noise schedule. We then train a denoiser $\bm{\epsilon}_{\theta}(\mathcal{J}_{t},t\mid\phi)$ that learns to predict the noise from the noised joints $\mathcal{J}_{t}=\{\mathbf{J}_{k,t}\}$ conditioned on the shape latent $\phi$, using the following loss:

### v2

@@ Abstract.
journalyear: 2025conference: Special Interest Group on Computer Graphics and Interactive Techniques Conference Conference Papers ; August 10–14, 2025; Vancouver, BC, Canadadoi: 10.1145/3721238.3730743isbn: 979-8-4007-1540-2/2025/08ccs: Computing methodologies Animation
@@ 1. Introduction
With rapid advances of immersive11footnotetext: Equal contribution, alphabetical order. Equal advising. media comes a growing demand for automated 3D content creation. 3D animation is essential nowadays across the entertainment industry, from game design and movie production to a wide range of AR/VR applications. Yet, producing realistic 3D animations remains one of the most labor-intensive tasks, often requiring enormous expert effort.
@@
The key insight is to leverage existing large-scale animated 3D assets handcrafted by expert artists. Large-scale 3D data has proven crucial in training high-quality 3D generation models (Shi et al., 2023; Liu et al., 2024). Notably, the Objaverse datasets (Deitke et al., 2022, 2024), aggregating over 10 million 3D assets from various public sources, have powered many recent state-of-the-art methods. However, prior efforts primarily use such assets for static 3D generation, neglecting the fact that many of these assets also come with valuable animation information skillfully crafted by the artists.
To train a learning-based model for 3D object rigging, we assemble a large-scale dataset of 230K rigged 3D assets curated from the Objaverse-XL Dataset (Deitke et al., 2024), each paired with artist-created rigging and skinning in a unified format, dubbed the Anymate Dataset. This dataset is 70 times larger than existing public rigging datasets, as summarized in Table 1, and contains a wide spectrum of 3D objects, ranging from humanoid and animal characters to animated everyday objects like furniture and machines.
@@ 3.2. Dataset Processing
We load these assets into Blender [^1] https://www.blender.org/ and extract the mesh, armature (bones), and skinning weights. For training efficiency, we resample 8,192 points on the mesh, and recompute their skinning weights using barycentric interpolation based on the original annotated skinning weights at the vertices. Finally, for each asset, we export the original mesh in a unified .obj format, as well as the resampled point clouds and the corresponding bone locations and skinning weights.
@@ 4.2.1. Architectures
A regression-based model is perhaps the most intuitive approach for joint prediction, which simply regresses a set of 3D coordinates from the extracted shape features and can be directly supervised by the ground-truth joints during training. However, there are two unique features that make our joint prediction task different from typical regression problems: (1) the number of target joints varies across instances, and (2) the order of the joints is immaterial at this stage [^2] Although an order could technically be established based on the skeleton, inconsistencies may arise across instances created by different artists.. To address these challenges, we design a transformer-based architecture (Jaegle et al., 2021; Vaswani et al., 2017), which is permutation-equivariant and has proven to be highly scalable, followed by a clustering step (Ester et al., 1996).
@@
where $\overline{\alpha}_{t}$ signifies the noise schedule. We then train a denoiser $\bm{\epsilon}_{\theta}(\mathcal{J}_{t},t\mid\phi)$ that learns to predict the noise from the noised joints $\mathcal{J}_{t}=\{\mathbf{J}_{k,t}\}$ conditioned on the shape latent $\phi$, using the loss: